OpenStudy (anonymous):

Solve for x: 3/x-2/3=5/6

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

3/x-2/3=5/6 3/x=5/6+2/3 3/x=9/6=3/2 x=2